OpenStudy (anonymous):

How to prove these two important properties of Frobenius norm : 1. For the identity matrix, ||I||F=Sqrt[n], whereas ||I||1=||I||infinity = 1.

OpenStudy (anonymous):

\[For the identity matrix, \left| \right|I \left| \right|_{F}=\sqrt{n}, whereas \left| \right|I \left| \right|_{1}=\left| \right|I \left| \right|_{\infty}=1.\]

OpenStudy (anonymous):

2. \[\left| \right|A \left| \right|_{F}^{2}=trace \left( A ^{T}A \right)\] where \[trace \left( A \right)\]is defined as the sum of the diagonal elements of A. That is, if \[A=\left( a _{ij} \right), then \] \[trace \left( A \right)=\sum_{i=1}^{n} a _{ii}\]

OpenStudy (anonymous):

I'm not even sure what the first one means.

OpenStudy (anonymous):

Doesn't say where \(F\) or \(n\) come from.

OpenStudy (anonymous):

F stand for Frebonius norm and n is for \[\mathbb{R} ^{n \times n}\]

OpenStudy (anonymous):

since I is identity matrix

OpenStudy (anonymous):

What is \[ \|I\|_1 \]Supposed to mean?

OpenStudy (anonymous):

its mean one norm form identity matrix.

OpenStudy (anonymous):

Okay can you actually give some definitions?

OpenStudy (anonymous):

You generally need a definition to prove properties.

OpenStudy (anonymous):

Okay..

OpenStudy (anonymous):

Given a matrix A and a vector norm \[\left| \right|.\left| \right|\] a non-negative number defined by \[\left| \right|A \left| \right|_{p}=\max_{x \neq 0}\frac{ \left| \right|A x \left| \right|_{p} }{ \left| \right|x \left| \right|_{p} }\]

OpenStudy (anonymous):

Okay now what is the \(\|_p\)?

OpenStudy (anonymous):

I only know about one norm

OpenStudy (anonymous):

owh, its refer to p-norm where p is a real number greater than or equal to 1.